OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−6, 4) and (2, 0).

OpenStudy (anonymous):

A line that has those points must have what slope?

OpenStudy (anonymous):

Use slope equation to find the slope first which canbe determined by. m=y2-y1/x2-x1 m=0-4/2-(-6) m=-4/8 m=-2 Now use the point slope form............